# ALMA detection of hydrogen cyanide (HCN) in the atmosphere of Saturn

### v4

@@ 3.5 Spatial distribution of HCN in Saturn
We create an integrated emission map of the HCN in Saturn using the task IMMOMENTS in CASA. Integrated emission maps are created by integrating the spectral data in the velocity range where the emission line of the HCN is detected. After the generation of the integrated emission map, we find that the emission line of HCN emit from the eastern and western limbs of Saturn. We overplot the integrated emission map of HCN over the 856.19 m continuum emission map of Saturn to determine the proper location of HCN emissions in Saturn. The integrated emission map and the overplotted map are shown in Figure 5. We observe the integrated map of HCN is resolved because the line emitting region is larger than synthesized beam of the integrated emission map.
@@ 3.6 Possible formation mechanism of HCN in the atmosphere of Saturn
Earlier, Tokunaga et al. (1981) claimed that the formation mechanism of HCN in the atmosphere of Saturn is similar to that of Jupiter. Tokunaga et al. (1981) stated that the photolysis of CH and NH is the most likely pathway for the formation of HCN in both Saturn and Jupiter. The laboratory experiment by Raulin et al. (1979) shows that HCN can be produced by photochemical reactions in a mixture of H, CH and NH under ultraviolet light. During the experiment, Raulin et al. (1979) observed that NH abundance is low in the stratosphere of Jupiter, and most of the HCN is produced by photolysis reactions in the lower stratosphere and upper troposphere of Jupiter. If this mechanism holds true in Saturn’s atmosphere, the abundance of HCN in Saturn is lower than that in Jupiter because Saturn’s atmosphere contains a much lower abundance of NH (Encrenaz et al., 1974; Tokunaga et al., 1981). Tokunaga et al. (1981) calculated the mixing ratio of HCN in Jupiter’s atmosphere to be (0.9–4.2)10, and we calculated the mixing ratio of HCN in Saturn’s atmosphere to be (2.90–6.19) ppb. This comparison indicates that the abundance of HCN in the atmosphere of Saturn is slightly less than the abundance of HCN in the atmosphere of Jupiter. This implies that the laboratory experiment of Raulin et al. (1979) is valid in the atmosphere of Saturn, i.e. the HCN in the atmosphere of Saturn is produced by the photolysis of CH and NH. Further photochemical modelling is needed to establish the most efficient pathways for the production of HCN in the atmosphere of Saturn.
@@ 4 Conclusion
In this letter, we analyzed the ALMA band 7 data for Saturn. The main results are as follows.
@@
The derived abundances of HCN in the western and eastern limbs were 6.19 ppb and 2.90 ppb, respectively. We also computed the TP and VMR retrival profiles using the PSG. After retrival analysis, we observed that the HCN is distributed up to 475 km in the stratosphere of Saturn. Similarly, from the TP profile, the temperature of HCN varied between 100 and 250 K.
We also discuss the possible formation mechanism of HCN towards Saturn and claim that the photolysis of CH and NH is responsible for the production of HCN.
@@ ACKNOWLEDGEMENTS
AM acknowledges the Swami Vivekananda Merit-cum-Means Scholarship (SVMCM) for financial support for this research. This paper makes use of the following data: ADS /JAO.ALMA#2017.1.00636.S. ALMA is a partnership of ESO, NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO, and NAOJ.

### v5

@@ 3.5 Spatial distribution of HCN in Saturn
We create an integrated emission map of the HCN in Saturn using the task IMMOMENTS in CASA. Integrated emission maps are created by integrating the spectral data cubes in the velocity ranges where the emission line of the HCN is detected. After the generation of the integrated emission map, we find that the emission line of HCN emits from the eastern and western limbs of Saturn. We overplot the integrated emission map of HCN over the 856.19 m continuum emission map of Saturn to determine the proper location of HCN emissions in Saturn. The integrated emission map and the overplotted map are shown in Figure 5. We observe that the integrated map of HCN is resolved because the line emitting region is larger than the synthesized beam of the integrated emission map.
@@ 3.6 Possible formation mechanism of HCN in the atmosphere of Saturn
Earlier, Tokunaga et al. (1981) claimed that the formation mechanism of HCN in the atmosphere of Saturn is similar to that of Jupiter. Tokunaga et al. (1981) stated that the photolysis of CH4 and NH3 is the most likely pathway for the formation of HCN in both Saturn and Jupiter. The laboratory experiment by Raulin et al. (1979) shows that HCN can be produced by photochemical reactions in a mixture of H2, CH4, and NH3 under ultraviolet light. During the experiment, Raulin et al. (1979) observed that NH3 abundance is low in the stratosphere of Jupiter, and most of the HCN is produced by photolysis reactions in the lower stratosphere and upper troposphere of Jupiter. If this mechanism holds true in Saturn’s atmosphere, the abundance of HCN in Saturn is lower than that in Jupiter because Saturn’s atmosphere contains a much lower abundance of NH3 (Encrenaz et al., 1974; Tokunaga et al., 1981). Tokunaga et al. (1981) calculated the mixing ratio of HCN in Jupiter’s atmosphere to be (0.9–4.2)10-9, and we calculated the mixing ratio of HCN in Saturn’s atmosphere to be (2.90–6.19) ppb. This comparison indicates that the abundance of HCN in the atmosphere of Saturn is slightly less than the abundance of HCN in the atmosphere of Jupiter. This implies that the laboratory experiment of Raulin et al. (1979) is valid in the atmosphere of Saturn, i.e., the HCN in the atmosphere of Saturn is produced by the photolysis of CH4 and NH3. Further photochemical modelling is needed to establish the most efficient pathways for the production of HCN in the atmosphere of Saturn.
@@ 4 Conclusion
In this paper, we analyzed the ALMA band 7 data for Saturn. The main results are as follows:
@@
The derived abundances of HCN in the western and eastern limbs are 6.19 ppb and 2.90 ppb, respectively. We also computed the TP and VMR retrieval profiles using the PSG. After retrieval analysis, we observed that the HCN is distributed up to 475 km in the stratosphere of Saturn. Similarly, from the TP profile, the temperature of HCN varied between 100 and 250 K.
We also discuss the possible formation mechanism of HCN towards Saturn and claim that the photolysis of CH4 and NH3 is responsible for the production of HCN.
@@ ACKNOWLEDGEMENTS
We thank the anonymous reviewers for her/his constructive comments that helped improve the manuscript. AM acknowledges the Swami Vivekananda Merit-cum-Means Scholarship (SVMCM) for financial support for this research. This paper makes use of the following data:
ADS /JAO.ALMA#2017.1.00636.S. ALMA is a partnership of ESO, NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO, and NAOJ.
Appendix